# MapAnything: Universal Feed-Forward Metric 3D Reconstruction map-anything.github.io

### v2

@@ 3.2 Factored Scene Representation Prediction
Once the multi-view transformer fuses information across different views and outputs the $N$-view patch tokens and scale token, MapAnything further decodes these tokens into factored quantities representing the metric 3D geometry. In particular, we use a DPT head [46] to decode the $N$-view patch tokens into $N$ dense per-view outputs, i.e., ray directions $R_{i}$ (normalized to unit length), up-to-scale ray depths $\tilde{D}_{i}$, masks $M_{i}$ representing non-ambiguous classes for depth, and world-frame point map confidence maps $C_{i}$. Furthermore, we also input the $N$-view patch tokens into an average pooling-based convolutional pose head [7] to predict the unit quaternions $Q_{i}$ and up-to-scale translations $\tilde{T}_{i}$. Finally, the scale token is passed through a 2-layer MLP with ReLU activations to predict the metric scaling factor. Since the metric scale of a scene can vary vastly, we exponentially scale the prediction to obtain the metric scaling factor $m$. As shown in Table˜5(a), we find that this decoupling of scale prediction is critical to achieving universal metric feed-forward inference. Finally, as mentioned earlier, these factored predictions can be used together to obtain the metric 3D reconstruction.
@@ 3.3 Training Universal Metric 3D Reconstruction
We train MapAnything end-to-end using multiple losses, depending on the available supervision. Since ray directions $R_{i}$ and pose quaternions $Q_{i}$ do not depend on scene scale, their losses are: $\mathcal{L}_{\text{rays}}\!=\!\sum_{i=1}^{N}\|\hat{R}_{i}\!-\!R_{i}\|$ and $\mathcal{L}_{\text{rot}}\!=\!\sum_{i=1}^{N}\min(\|\hat{Q}_{i}\!-\!Q_{i}\|,\|{-}\hat{Q}_{i}\!-\!Q_{i}\|)$. This accounts for the two-to-one mapping of unit quaternions and the regression loss is similar to a geodesic angular distance.
@@
For the factored predictions, we find up-weighting the global pointmap loss and down-weighting the mask loss to be beneficial. For all the regression losses, we use an adaptive robust loss [4] with $c=0.05$ and $\alpha=0.5$ to help with inliers.
@@
To enable one-shot training of a universal model that can support various input configurations, we provide additional geometric inputs to the model with varying selection probabilities during training. Specifically, we use an overall geometric input probability of 0.9, where each individual factorization, i.e., ray directions, ray depth, and pose, has an input probability of 0.5 each. Whenever depth is selected as input, there is an equal probability of providing dense depth or 90% randomly sparsified depth. For robustness and flexibility in terms of which views have geometric information available as input, we use a per-view input probability of 0.95 and do not provide metric scale factors as input for metric-scale ground-truth datasets with a probability of 0.05. We provide further details regarding the training setup in the supplement.
@@
We train MapAnything on 13 high-quality datasets (see Table˜1) with diversity across indoor, outdoor, and in-the-wild scenes. For ScanNet++ v2 and TartanAirV2-WB, we split the scenes into a training, validation, and a held-out test set, while other datasets are split into training and validation. While MPSD is originally a monocular metric depth dataset, we acquire the pose and camera information to enable a real-world multi-view metric scale dataset with  72K scenes. We will be open-sourcing this MPSD metadata to enable future research. We release two pretrained models: one licensed under Apache 2.0 trained on six datasets, and one licensed under CC BY-NC 4.0 trained on an additional seven datasets (see Table˜1). We provide comparisons between both variants in the supplementary.
Figure 5: MapAnything shows state-of-the-art dense multi-view reconstruction for number of input views varying from 2 to 100 and under different input configurations. We report the absolute relative error (rel), the inlier ratio at a relative threshold of $1.03\%$ ($\tau$), the average aligned trajectory error (ATERMSEATE~RMSE), the area under the curve at an error threshold of 5° ($AUC@5$), and the average angular error (errerr) in degrees (°), averaged over ETH3D, ScanNet++ v2 & TAv2. We don’t report performance for baselines when the inference runs out of GPU memory. We provide results for the exhaustive input configurations of MapAnything in the supplement.
@@
Figure 6: MapAnything provides high-fidelity dense geometric reconstructions across varying domains and number of views. Here we are showcasing its capabilities only using images as input. It also works well on monocular and art images despite not being trained for it.
@@
MapAnything 0.01 0.02 92.1 0.00 93.2 0.01 91.6 0.14
4 Benchmarking & Results
In this section, we benchmark MapAnything across a wide suite of 3D vision tasks. For each task, we compare against expert baselines specifically designed or trained for the task. We perform all the experiments with a constant seed.
Multi-View Dense Reconstruction:
We benchmark the performance of pointmaps, pose, depth & ray directions estimation on an undistorted version of ETH3D [53], ScanNet++ v2 [82], and TartanAirV2-WB [73, 87], where, for each test scene, we randomly sample up to $N$ views that form a single connected component graph based on the pre-computed pair-wise covisibility of all images in the scene (this prevents disjoint sets of images as input). Figure˜5 shows that MapAnything provides state-of-the-art dense multi-view reconstruction performance over other baselines using only image input, including VGGT [66]. Beyond the performance using only images as input, we show that MapAnything can leverage additional auxiliary geometric inputs for feed-forward inference to further increase reconstruction performance by a significant factor. Furthermore, we find MapAnything is better than the bundle adjustment (BA) variant of the two-view baseline, Pow3R [25], which is also designed to leverage scene priors. In Figure˜3, we illustrate how the auxiliary geometric inputs improve MapAnything. We also find that the reconstruction outputs from MapAnything (using only images as input) display high fidelity, as shown in Figure˜4.
Two-View Dense Reconstruction:
We benchmark sparse-view reconstruction and image matching performance against state-of-the-art feed-forward baselines in Table˜2. MapAnything achieves state-of-the-art performance using only images as input. With additional input modalities, MapAnything significantly outperforms both image-only baselines and Pow3R [25], the only other two-view feed-forward method that uses scene or camera priors.
@@ 4 Benchmarking & Results
We benchmark the single-view calibration performance of MapAnything and other expert calibration baselines on randomly sampled frames from the test scenes of undistorted ETH3D [53], ScanNet++ v2 [82], and TartanAirV2 [73]. To test non-centered principal points, we randomly crop frames with aspect ratios from 3:1 to 1:2. Despite not being trained specifically for single images, Table˜3 shows that MapAnything achieves state-of-the-art performance for perspective calibration. This demonstrates MapAnything’s effectiveness in modeling generic central camera systems and its potential to generalize to wide-angle models like fisheye with appropriate training.
Table 3: MapAnything showcases state-of-the-art single image calibration. Note that MapAnything has not been trained specifically for single image inputs. We report the average angular error (errerr) in degrees (°). Best results are indicated in bold.
@@
MapAnything 1.18 1.70 0.36 1.47
Table 4: MapAnything showcases versatile metric depth estimation under different input configurations on the Robust-MVD Benchmark [50]. Note that MapAnything has not been trained for single image inputs. We report the absolute relative error (rel) and the inlier ratio at a relative threshold of 1.03% ($\tau$). The best result for each group is in bold; gray text indicates results where the evaluation dataset is in the training distribution [23]. We provide the full version with alignment results in the supplement.
KITTI [12] ScanNet [8]
@@
MapAnything 04.23 56.4 017.31 11.9
@@
In Table˜4, we benchmark MapAnything against expert models for single-view and multi-view depth estimation across various inputs. Although not trained for single-view metric depth, MapAnything achieves state-of-the-art or comparable performance. For multi-view metric depth estimation using images only, MapAnything outperforms MASt3R-BA [30] & MUSt3R [6]. With auxiliary inputs like camera calibration and poses, MapAnything’s performance improves and it delivers competitive results to task-specific expert models. In comparison to baselines such as MoGe-2 [71] and MVSA [24], we find the metric scale estimation/pass through on ScanNet to be sub-optimal and believe this might be due to low benchmarking dataset quality [69, 24]. As indicated in Table˜S.2, we observe strong depth estimation performance on ScanNet when using scale alignment.
@@
As showcased in Table˜5(a), the factored representation of the scene as a multi-view set of rays, depth & pose (RDP) along with the metric scale is a key enabler for strong reconstruction performance while using images and optionally additional geometric inputs. In Table˜5(b), we find that our input probability based training is efficient in training one universal model for various tasks and input configurations, where the performance of the universally trained model is equivalent to various bespoke models trained for the specific input configurations.
@@
Qualitative Examples:
We showcase visualizations of diverse MapAnything reconstructions in Figure˜6. While MapAnything can support a flexible set of inputs, MapAnything already showcases robust dense reconstruction using only images as input. In particular, we show results across varying number of captured views and from different domains such as indoor, landscape, art, object-centric, and off-road.
@@ 5 Limitations
While MapAnything makes significant strides towards a universal multi-modal backbone for in-the-wild metric-scale 3D reconstruction, several limitations and future directions remain: (a) MapAnything does not explicitly account for the noise or uncertainty in geometric inputs. (b) Although this is not currently supported, the architecture can be easily extended to handle tasks where images are not available for all input views. For example, in novel view synthesis, the target views for rendering will only have cameras available as input. (c) While the design of MapAnything supports iterative inference, it is yet to be explored how effective scaling of test-time compute would be for 3D reconstruction (this ties into effectively handling noise in the inputs). (d) Multi-modal features are currently fused before input; exploring ways for efficient direct input of different modalities to the transformer could be interesting.
Beyond multi-task capabilities, scalability is limited by the one-to-one mapping between input pixels and the output scene representation. We believe that significant work remains in effectively representing scenes in memory and decoding them as required, especially for large scenes. Furthermore, our current scene parameterization does not capture dynamic motion, or scene flow, which are promising areas.
@@ 6 Conclusion
MapAnything is the first universal transformer-based backbone that directly regresses metric 3D geometry and camera poses from flexible inputs – including images, camera intrinsics, poses, depth maps, or partial reconstructions – in a single pass. By using a factored representation of multi-view geometry (depth maps, ray maps, poses, and a global scale factor), MapAnything unifies local estimates into a global metric frame. With standardized supervision across varied datasets and augmentations, MapAnything handles multiple tasks like uncalibrated structure-from-motion, calibrated multi-view stereo, monocular depth estimation, camera localization, depth completion, and more without task-specific tuning. Extensive experiments show that it surpasses or matches specialist models while enabling efficient joint training. Future extensions to dynamic scenes, uncertainty quantification, and scene understanding promise to further generalize MapAnything’s capabilities and robustness, paving the way toward a truly universal 3D reconstruction backbone.
Acknowledgments
@@
MapAnything: Universal Feed-Forward Metric 3D Reconstruction
Supplementary Material

### v3

@@ 3.2 Factored Scene Representation Prediction
Once the multi-view transformer fuses information across different views and outputs the $N$-view patch tokens and scale token, MapAnything further decodes these tokens into factored quantities representing the metric 3D geometry. In particular, we use a DPT head [46] to decode the $N$-view patch tokens into $N$ dense per-view outputs, i.e., ray directions $R_{i}$ (normalized to unit length), up-to-scale ray depths $\tilde{D}_{i}$, masks $M_{i}$ representing non-ambiguous classes for depth, and world-frame pointmap confidence maps $C_{i}$. Furthermore, we input the $N$-view patch tokens into an average pooling-based convolutional pose head [7] to predict the unit quaternions $Q_{i}$ and up-to-scale translations $\tilde{T}_{i}$. Finally, the scale token is passed through a 2-layer MLP with ReLU activations to predict the metric scaling factor. Since the metric scale of a scene can vary vastly, we exponentially scale the prediction to obtain the metric scaling factor $m$. As shown in Table˜5(a), we find that this decoupling of scale prediction is critical to achieving universal metric feed-forward inference. Finally, as mentioned earlier, these factored predictions can be used together to obtain the metric 3D reconstruction.
@@ 3.3 Training Universal Metric 3D Reconstruction
We train MapAnything end-to-end using multiple losses depending on the available supervision. Since ray directions $R_{i}$ and pose quaternions $Q_{i}$ do not depend on scene scale, their losses are: $\mathcal{L}_{\text{rays}}\!=\!\sum_{i=1}^{N}\|\hat{R}_{i}\!-\!R_{i}\|$ and $\mathcal{L}_{\text{rot}}\!=\!\sum_{i=1}^{N}\min(\|\hat{Q}_{i}\!-\!Q_{i}\|,\|{-}\hat{Q}_{i}\!-\!Q_{i}\|)$. This accounts for the two-to-one mapping of unit quaternions, and the regression loss is similar to a geodesic angular distance.
@@
For the factored predictions, we find that up-weighting the global pointmap loss and down-weighting the mask loss is beneficial. For all the regression losses, we use an adaptive robust loss [4] (with parameters $c=0.05$ and $\alpha=0.5$) to help with robustness to outliers.
@@
To enable one-shot training of a universal model that supports various input configurations, we provide additional geometric inputs to the model with varying selection probabilities during training. Specifically, we use an overall geometric input probability of 0.9, where each individual factorization, i.e., ray directions, ray depth, and pose, has an input probability of 0.5 each. Whenever depth is selected as input, there is an equal probability of providing dense depth or 90% randomly sparsified depth. For robustness and flexibility in terms of which views have geometric information available as input, we use a per-view input probability of 0.95 and do not provide metric scale factors as input for metric-scale ground-truth datasets with a probability of 0.05. We provide further details regarding the training setup in the supplement.
@@
We train MapAnything on 13 high-quality datasets (see Table˜1) with diversity across indoor, outdoor, and in-the-wild scenes. For ScanNet++ v2 and TartanAirV2-WB, we split the scenes into a training, validation, and a held-out test set, while other datasets are split into training and validation. While MPSD is originally a monocular metric depth dataset, we acquire the pose and camera information to enable a real-world multi-view metric scale dataset with $\sim$72K scenes. We open-sourced this MPSD metadata to enable future research. We release two pretrained models: one licensed under Apache 2.0 trained on six datasets, and one licensed under CC BY-NC 4.0 trained on an additional seven datasets (see Table˜1). We provide comparisons between both variants in the supplementary.
@@
4 Benchmarking & Results
In this section, we benchmark MapAnything across a wide suite of 3D vision tasks. For each task, we compare against expert baselines specifically designed or trained for the task. We perform all experiments with a constant seed.
Multi-View Dense Reconstruction:
We benchmark the performance of pointmaps, pose, depth & ray direction estimation on an undistorted version of ETH3D [54], ScanNet++ v2 [81], and TartanAirV2-WB [73, 86], where, for each test scene, we randomly sample up to $N$ views that form a single connected component graph based on the pre-computed pairwise covisibility of all images in the scene (this prevents disjoint sets of images as input). Figure˜4 shows that MapAnything provides state-of-the-art dense multi-view reconstruction performance over other baselines using only image input, including VGGT [67]. Beyond the performance using only images as input, we show that MapAnything can leverage additional auxiliary geometric inputs for feed-forward inference to further increase reconstruction performance by a significant factor. Furthermore, we find that MapAnything is better than the bundle adjustment (BA) variant of the two-view baseline, Pow3R [23], which is also designed to leverage scene priors. We also find that reconstruction outputs from MapAnything (using only images as input) display high fidelity, as shown in Figure˜3.
Figure 4: MapAnything shows state-of-the-art dense multi-view reconstruction for input views ranging from 2 to 100 and under different input configurations. We report the absolute relative error (rel), the inlier ratio at a relative threshold of $1.03\%$ ($\tau$), the average aligned trajectory error (ATERMSEATE~RMSE), the area under the curve at an error threshold of 5° ($AUC@5$), and the average angular error (errerr) in degrees (°), averaged over ETH3D, ScanNet++ v2 & TAv2. We do not report performance for baselines when the inference runs out of GPU memory. We provide results for individual datasets & the exhaustive input configurations of MapAnything in the supplement.
Two-View Dense Reconstruction:
We benchmark sparse-view reconstruction and image-matching performance against state-of-the-art feed-forward baselines in Table˜2. MapAnything achieves state-of-the-art performance using only images as input. With additional input modalities, MapAnything significantly outperforms both image-only baselines and Pow3R [23], the only other two-view feed-forward method that uses scene or camera priors.
@@ 4 Benchmarking & Results
MapAnything 0.01 0.02 82.0 0.00 94.8 0.02 81.5 0.16
@@
We benchmark the single-view calibration performance of MapAnything and other expert calibration baselines on randomly sampled frames from the test scenes of undistorted ETH3D [54], ScanNet++ v2 [81], and TartanAirV2 [73]. To test non-centered principal points, we randomly crop frames with aspect ratios from 3:1 to 1:2. Despite not being trained specifically on single images, Table˜3 shows that MapAnything achieves state-of-the-art performance for perspective calibration. This demonstrates MapAnything’s effectiveness in modeling generic central camera systems and its potential to generalize to wide-angle models like fisheye with appropriate training.
Table 3: MapAnything shows state-of-the-art single-image calibration. Note that MapAnything has not been trained specifically for single-image inputs. We report the average angular error (errerr) in degrees (°). Best results are indicated in bold.
@@
MapAnything 1.06 1.33 0.39 1.47
Table 4: MapAnything shows versatile metric depth estimation under different input configurations on the Robust-MVD Benchmark [51]. Note that MapAnything has not been trained for single-image inputs. We report the absolute relative error (rel) and the inlier ratio at a relative threshold of 1.03% ($\tau$). The best result for each group is in bold; gray text indicates results where the evaluation dataset is in the training distribution [22].
KITTI ScanNet
@@
MapAnything 04.63 51.6 005.58 48.1
c) Single-View w/ Alignment
MoGe [70] 05.12 46.2 003.59 65.3
MoGe-2 [71] 04.82 47.9 003.77 63.1
VGGT [67] 07.50 33.0 003.33 70.8
$\pi^{3}$ [74] 06.00 40.1 002.90 73.9
MapAnything 06.12 42.2 004.95 55.6
Depth Pro [5] 06.10 39.6 004.30 58.4
DAV2 [79] 06.60 38.6 004.00 58.6
Metric3DV2 [18] 05.10 44.1 002.40 78.3
UniDepthV2 [45] 04.00 55.3 002.10 82.6
MapAnything 06.15 41.6 004.77 57.1
d) Multi-View w/ Alignment
MAST3R [29] 03.30 67.7 004.30 64.0
MUSt3R [6] 04.47 56.7 003.22 69.2
VGGT [67] 04.60 53.0 002.34 80.6
$\pi^{3}$ [74] 03.09 69.5 001.98 83.6
MapAnything 04.04 60.3 003.47 67.0
DeMoN [61] 15.50 15.2 012.00 21.0
DeepV2D KITTI [57] 03.10 74.9 023.70 11.1
DeepV2D ScanNet [57] 10.00 36.2 004.40 54.8
MapAnything 03.97 61.2 003.34 68.5
@@
In Table˜4, we benchmark MapAnything against specialized models for single-view and multi-view depth estimation across various inputs. In the RMVD benchmark, note that we don’t use ETH3D due to the distortion issue mentioned in MVSA [22] and DTU & Tanks and Temples since they are not metric. Although not trained specifically for single-view metric depth, MapAnything achieves state-of-the-art or comparable performance. For multi-view metric depth estimation using images only, MapAnything outperforms MASt3R-BA [29] & MUSt3R [6]. With auxiliary inputs like camera calibration and poses, MapAnything’s performance improves and it delivers competitive results compared to task-specific specialized models. In comparison to baselines such as MoGe-2 [71] and MVSA [22], we find the metric scale estimation on ScanNet to be sub-optimal and believe this is likely due to lower benchmark dataset quality [70, 22]. As indicated in Table˜4, we observe strong depth estimation performance on ScanNet when using median scale alignment.
@@
As shown in Table˜5(a), the factored representation of the scene as a multi-view set of rays, depth & pose (RDP) along with the metric scale is a key enabler for strong reconstruction performance while using images and optionally additional geometric inputs. In Table˜5(b), we find that our input probability-based training is efficient in training one universal model for various tasks and input configurations, where the performance of the universally trained model is equivalent to various bespoke models trained for specific input configurations.
@@ 5 Limitations
While MapAnything makes significant strides towards a universal multi-modal backbone for in-the-wild metric-scale 3D reconstruction, several limitations and future directions remain: (a) MapAnything does not explicitly account for noise or uncertainty in geometric inputs. (b) Although this is not currently supported, the architecture can be easily extended to handle tasks where images are not available for all input views. For example, in novel view synthesis, the target views for rendering will only have cameras available as input. (c) While the design of MapAnything supports iterative inference, it remains to be explored how effective scaling of test-time compute would be for 3D reconstruction (this ties into effectively handling noise in the inputs). (d) Multi-modal features are currently fused before being input; exploring more efficient ways to directly input different modalities to the transformer could be interesting.
Beyond multi-task capabilities, scalability is currently limited by the one-to-one mapping between input pixels and the output scene representation. We believe that significant work remains in effectively representing scenes in memory and decoding them as required, especially for large scenes. Our current scene parameterization does not capture dynamic motion or scene flow [26], which are promising areas.
@@ 6 Conclusion
MapAnything is the first universal transformer-based backbone that directly regresses metric 3D geometry and camera poses from flexible inputs – including images, camera intrinsics, poses, depth maps, or partial reconstructions – in a single pass. By using a factored representation of multi-view geometry (depth maps, ray maps, poses, and a global scale factor), MapAnything unifies local estimates into a global metric frame. MapAnything handles multiple tasks like uncalibrated structure-from-motion, calibrated multi-view stereo, monocular depth estimation, camera localization, depth completion, and more without task-specific tuning. Extensive experiments show that it surpasses or matches specialist models while enabling efficient joint training. Future extensions to dynamic scenes, uncertainty quantification, and scene understanding promise to further generalize MapAnything’s capabilities and robustness, paving the way toward a truly universal 3D reconstruction backbone.
MapAnything: Universal Feed-Forward Metric 3D Reconstruction
Supplementary Material
Appendix A Acknowledgments
@@
Appendix B Implementation Details
[…]
Appendix C Additional Evaluation
Speed & Memory Profiling:
We present profiling results for MapAnything and other concurrent models in Figure˜S.1.
Flexibility of Input Configurations:
We present a representative set of input configurations for MapAnything in Table˜S.2, where the performance of MapAnything improves as more modalities are provided. The universal training of MapAnything with varying selection probabilities for geometric inputs as 6 factors (as described in Section˜3.3) enables support for 64 exhaustive input combinations. While we primarily benchmark cases where the input modality is available for all views, MapAnything can also support optional geometric inputs for a subset of the input views.
Comparison of MapAnything Variants:
We provide a comparison between different variants of MapAnything in Figure˜S.2, where the performance of VGGT [67] is provided as a baseline. First, we show that our two-stage training with covisibility-based view sampling is very effective, where the MapAnything model trained for up to 4 views as input already shows strong generalization to a significantly higher number of views. We also compare the performance of our different open-source models, where, as described in Table˜1, the Apache model is trained on 6 datasets, while the CC-BY-NC one is trained on 13. While we observe a decrease in performance, the Apache variant is still competitive with the VGGT baseline and its performance further improves as additional geometric inputs are provided. We also provide a comparison between the various versions of the publicly released models in Figure˜S.3.
Design Choices:
As shown in Table˜S.2, log scaling and alternating attention are key for strong performance when evaluating with 50 views, well beyond the 4 used in training.
Qualitative Examples:
Figure˜S.4 and Figure˜S.5 showcase the versatility of MapAnything.
Comparison with Concurrent Models:
Figure˜S.6 showcases the average results, while Figure˜S.7 and Figure˜S.8 show the performance for each individual dataset.
Figure S.1: MapAnything shows the best speed and memory usage profile in comparison to latest concurrent state-of-the-art multi-view feed-forward reconstruction models. We profile all the models on a H200-140GB GPU using appropriate floating point precision autocast. Upon profiling the naive inference call of MapAnything and comparing to concurrent methods, we notice a memory bottleneck in the dense per-pixel regression across views, since our implementation of the dense head is closer to the original DPT [46]. To alleviate this issue, i.e., in MapAnything (Mem Efficient), we run the per-view decodings in a mini-batched loop, where in the above profiling the mini-batch size is 1. We find that this looping leads to negligible trade-off in speed while significantly reducing the memory usage.
Figure S.2: The Apache and first stage (up to 4 views) training variants of MapAnything show strong dense multi-view reconstruction for input views ranging from 2 to 100 and under different input configurations. We report the absolute relative error (rel), the inlier ratio at a relative threshold of $1.03\%$ ($\tau$), the average aligned trajectory error (ATERMSEATE~RMSE), the area under the curve at an error threshold of 5° ($AUC@5$), and the average angular error (errerr) in degrees (°), averaged over ETH3D, ScanNet++ v2 & TAv2.
Figure S.3: Dense multi-view reconstruction benchmarking comparing the performance of the latest checkpoint as of January 20th 2026 (MapAnything) to the one released in September 2025 (MapAnything-Prev). The primary difference between the two variants (latest & previous) is: (a) Encoder: First 24 layers of 1536-dim DINOv2 ViT-G (latest) vs All 24 layers of 1024-dim DINOv2 ViT-L (previous), (b) Multi-View Transformer: Last 16 layers of 1536-dim DINOv2 ViT-G (latest) vs 24 layers of a randomly initialized 736-dim ViT-B size transformer. Both across large-scale trainings and ablation settings, we find that the DINOv2 initialization for the multi-view transformer helps significantly with convergence speed and final performance. In the above plots, we report the absolute relative error (rel), the inlier ratio at a relative threshold of $1.03\%$ ($\tau$), the average aligned trajectory error (ATERMSEATE~RMSE), the area under the curve at an error threshold of 5° ($AUC@5$), and the average angular error (errerr) in degrees (°), averaged over ETH3D, ScanNet++ v2 & TAv2.
Figure S.4: MapAnything provides high-fidelity dense geometric reconstructions across varying domains and number of views. While MapAnything can support a flexible set of inputs, here we are showcasing its capabilities only using images as input. In particular, we show results across a varying number of captured views and from different domains such as indoor, landscape, art, object-centric, and off-road. It also works well on monocular and art images despite not being trained for it.